Equivalent Units of Production

The Converting Department of Hopkinsville Company had 1,120 units in work in process at the beginning of the period, which were 25% complete. During the period, 23,600 units were completed and transferred to the Packing Department. There were 1,240 units in process at the end of the period, which were 25% complete. Direct materials are placed into the process at the beginning of production.

Determine the number of equivalent units of production with respect to direct materials and conversion costs. If an amount is zero, enter in "0".

Hopkinsville Company
Number of Equivalent Units of Production
Whole Units Direct Materials Equivalent Units Conversion Equivalent Units
Inventory in process, beginning
Started and completed
Transferred to Packing Department
Inventory in process, ending
Total
Answer #1
Hopkinsville Company
Number of Equivalent Units of Production
Whole Units Direct Materials Equivalent Units Conversion Equivalent Units
Inventory in process, beginning 1120 0 840 =1120*(1-25%)
Started and completed 22480 22480 22480 =23600-1120
Transferred to Packing Department 23600 22480 23320
Inventory in process, ending 1240 1240 310 =1240*25%
Total 24840 23720 23630